# Rethinking Popularity Bias in Collaborative Filtering via Analytical Vector Decomposition

### v5

@@ 1. Introduction
Collaborative filtering (CF), which models user preferences based on historical interactions, serves as the cornerstone of modern recommender systems (DisentanglingUserZheng2021; ImprovingRecommendationChen2023; ASurveyonAccuracyWu2023; wang2018confidence; wang2020setrank). Embedding-based models, particularly those leveraging graph neural networks like LightGCN (LightGCNHe2020), have achieved great success by learning user and item representations in a latent space (GraphNeuralWu2023; wang2021variable; wang2021personalized). These models are almost universally trained with the Bayesian Pairwise Ranking (BPR) loss, a de facto standard for implicit feedback datasets. However, BPR optimization exhibits a fundamental geometric flaw: it systematically distorts the embedding space to favor popular items, creating a popularity bias that undermines personalization (BiasandDebiasChen2023; PopularityBiasZhu2021). This bias not only reduces recommendation personalization and diversity (AutoDebiasChen2021; ItemSideJiang2024; FairCFShao2022), but also perpetuates a Matthew effect where popular items dominate while niche content is systematically overlooked (ASurveyontheFairnesWang2023; zhu2024graph; wang2025face).
Traditional approaches in recommender systems typically view popularity bias as an external confounding factor that interferes with recommendation quality, rather than recognizing it as an intrinsic consequence of the optimization process itself. These methods employ causal inference or additional bias modeling techniques to separately address bias effects without integrating deeply with the embedding learning mechanism. For example, re-weighting methods, such as inverse propensity scoring (IPS), attempt to balance training data by down-weighting popular item interactions (CounterfactualReasoningBottou2013; AutoDebiasChen2021). Regularization approaches penalize correlations between recommendation scores and item popularity (IncorporatingDiversityJacek2016; ControllingPopularityAbdollahpouri2017). Recently, causal methods model and remove confounding effects (DeconfoundedRecommendationWang2021; CausalInterventionZhang2021) through causal analyses. However, these macroscopic approaches address symptoms rather than root causes (ModelAgnosticWei2021). They fail to answer a fundamental question: how does BPR optimization geometrically distort the latent representation space to systematically favor popular items? Understanding this geometric mechanism is crucial for developing principled solutions that address bias at its source.
@@ 2.2.1. Matrix Factorization (MF)
Matrix Factorization (BPRRendle2009) is a foundational collaborative filtering model. It directly learns the user and item embeddings, $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$, from the interaction data. The preference score $\hat{y}_{ui}$ that user $u$ is predicted to have for item $i$ is calculated as the inner product of their respective embeddings:
@@ 2.2.2. LightGCN
LightGCN (LightGCNHe2020) is a state-of-the-art GNN-based model that simplifies the design of traditional Graph Convolutional Networks (GCNs) for recommendation. LightGCN begins with initial, learnable embeddings $\mathbf{E}^{(0)}=[\mathbf{e}_{u_{1}}^{(0)},...,\mathbf{e}_{i_{1}}^{(0)},...]^{T}$ and refines them through a linear propagation mechanism on the user-item interaction graph. At each layer $k$, the embeddings are aggregated from their neighbors as follows:
@@ 2.3. Bayesian Pairwise Ranking (BPR) Loss
The Bayesian Pairwise Ranking (BPR) loss is the standard optimization objective for models trained on implicit feedback datasets (BPRRendle2009). It is built on a pairwise learning assumption: for a given user, an observed item should be ranked higher than all unobserved items. To operationalize this, BPR constructs training instances as triplets $(u,i,j)$, where $u$ is a user, $i$ is an item they have interacted with (a positive item, $(u,i)\in\mathcal{R}$), and $j$ is an item they have not interacted with (a negative item, $(u,j)\notin\mathcal{R}$). The objective is to maximize the score difference between the positive and negative items, $\hat{y}_{ui}-\hat{y}_{uj}$. The BPR loss function is formulated as:
@@ 5.1.2. Baselines
  • Backbone Models: We show the general applicability of DDC by applying it to five representative models: MF (BPRRendle2009), LightGCN (LightGCNHe2020), DGCF (DisentangledGraphWang2020), NCL (ImprovingGraphLin2022), and LightCCF (UnveilingContrastiveZhang2025).
  • Debiasing Methods: We compare DDC with seven state-of-the-art debiasing methods: IPS (UnbiasedLearningJoachims2017), DICE (DisentanglingUserZheng2021), MACR (ModelAgnosticWei2021), PC (PopularityOpportunityZhu2021), PAAC (PopularityAwareCai2024), DCCL (DCCL), and TPAB (TPAB). For a fair comparison, these methods are applied to two backbone models, MF and LightGCN.
@@ 5.1.4. Implementation Details
To ensure a fair comparison, we set the training batch size to 8192 and implement all methods under the same framework, RecBole (recbole[1.0]; recbole[2.0]; recbole[1.2.1]). For all models, including backbones and debiasing baselines, we performed a grid search to find optimal hyperparameters. Crucially, we did not set an upper limit on epochs; all models were trained until convergence, determined by an early stopping strategy where training terminates if validation performance (MRR@10) does not improve for 50 consecutive epochs. The model achieving the best validation performance was then evaluated on the test set. The embedding dimension is set to $d=64$. For our DDC method, the only significant hyperparameter is the proportion of a user’s interacted items used to construct their personalized preference direction $\mathbf{e}_{pref}$. While careful tuning of this proportion can yield further gains, for the main experiments in Tables 2 and 3, we uniformly set this value to 30% to demonstrate robust performance without extensive hyperparameter tuning.
@@ 6. Related Work
Macroscopic Debiasing Strategies. Early approaches to mitigating popularity bias treat reommendation models as black boxes, intervening at the data or objective function level. For example, reweighting methods, like inversion propensity scoring (IPS) (CounterfactualReasoningBottou2013; AutoDebiasChen2021; PopularityOpportunityZhu2021), down-weight popular item interactions to amplify signals from niche items, but can suffer from high variance and oversimplification (UnbiasedLearningJoachims2017; RecommendationsasTreatmentsSchnabel2016). Regularization techniques add penalties to the objective function to discourage correlation with item popularity, often at the cost of overall recommendation accuracy (IncorporatingDiversityJacek2016; ControllingPopularityAbdollahpouri2017; PopularityOpportunityZhu2021; 10.1145/3711896.3737068). Critically, these macroscopic strategies address the symptoms of popularity bias without dissecting its underlying structural impact on the model’s learned representations.
Causal and Disentangled Approaches. More principled solutions leverage causal inference and representation disentanglement. Causal approaches use graphs to model and remove popularity’s confounding influence, often via counterfactual inference (DeconfoundedRecommendationWang2021; CausalInterventionZhang2021; CausalEmbeddingsBonner2018). Similarly, disentanglement methods aim to learn separate latent factors for genuine interest versus conformity to popular trends (DisentanglingUserZheng2021), with the goal of isolating a ‘pure’ preference representation for prediction. While sophisticated, these methods often rely on strong causal assumptions or complex training schemes (ModelAgnosticWei2021). Crucially, they hypothesize a popularity component but fail to explain how it is structurally and geometrically encoded by standard optimization like BPR—a foundational gap our work directly addresses.
Representation Space Optimization. Contrastive learning (CL) has emerged as a powerful paradigm to improve embedding quality in GNN-based models (SelfSupervisedWu2021; AreGraphAugmentationsYu2022). CL-based methods like SGL (SelfSupervisedWu2021) and SimGCL (AreGraphAugmentationsYu2022) employ a contrastive loss to foster a more uniform distribution of embeddings, which can indirectly alleviate popularity bias by preventing popular items from dominating the latent space (UnderstandingContrastiveWang2020; TowardsRepresentationWang2022). However, these approaches pursue uniformity as a general objective for representation quality; they do not directly address the specific, systematic geometric distortion that popularity imprints on the embedding space. Therefore, while beneficial, this pursuit of general uniformity is an indirect and incomplete solution to the specific geometric problem of popularity bias. Our work differs by identifying the principal “popularity direction” and then applying a targeted, asymmetric correction, fundamentally realigning the space to disentangle preference from popularity.
@@ Acknowledgements.
Table 6. Table of Notations
Symbol Description
$\mathcal{U}$ The finite set of all users, $u\in\mathcal{U}$.
$\mathcal{I}$ The finite set of all items, $i\in\mathcal{I}$.
$\mathcal{S}$ The set of observed user-item interactions, $\mathcal{S}\subseteq\mathcal{U}\times\mathcal{I}$.
$\mathcal{I}_{u}^{+}$ The set of items user $u$ has interacted with, $\{i\in\mathcal{I}\mid(u,i)\in\mathcal{S}\}$.
$\mathcal{U}_{i}^{+}$ The set of users who have interacted with item $i$, $\{u\in\mathcal{U}\mid(u,i)\in\mathcal{S}\}$.
$d\in\mathbb{N}$ The dimensionality of the embedding space.
$\mathbf{e}_{u}\in\mathbb{R}^{d}$ The embedding vector for user $u$.
$\mathbf{e}_{i}\in\mathbb{R}^{d}$ The embedding vector for item $i$.
$\hat{r}_{ui}$ The predicted score for a user-item pair $(u,i)$.
$\sigma(x)$ The logistic sigmoid function, $1/(1+e^{-x})$.

### v6

@@ 1. Introduction
Collaborative filtering (CF), which models user preferences based on historical interactions, serves as the cornerstone of modern recommender systems (Zheng et al., 2021; Chen et al., 2023b; Wu et al., 2023; Wang et al., 2018, 2020a). Embedding-based models, particularly those leveraging graph neural networks like LightGCN (He et al., 2020), have achieved great success by learning user and item representations in a latent space (Wu et al., 2022; Wang et al., 2021a, b). These models are almost universally trained with the Bayesian Pairwise Ranking (BPR) loss, a de facto standard for implicit feedback datasets. However, BPR optimization exhibits a fundamental geometric flaw: it systematically distorts the embedding space to favor popular items, creating a popularity bias that undermines personalization (Chen et al., 2023a; Zhu et al., 2021a). This bias not only reduces recommendation personalization and diversity (Chen et al., 2021; Jiang et al., 2024; Shao et al., 2022), but also perpetuates a Matthew effect where popular items dominate while niche content is systematically overlooked (Wang et al., 2023; Zhu et al., 2024; Wang et al., 2025).
Traditional approaches in recommender systems typically view popularity bias as an external confounding factor that interferes with recommendation quality, rather than recognizing it as an intrinsic consequence of the optimization process itself. These methods employ causal inference or additional bias modeling techniques to separately address bias effects without integrating deeply with the embedding learning mechanism. For example, re-weighting methods, such as inverse propensity scoring (IPS), attempt to balance training data by down-weighting popular item interactions (Bottou et al., 2013; Chen et al., 2021). Regularization approaches penalize correlations between recommendation scores and item popularity (Wasilewski and Hurley, 2016; Abdollahpouri et al., 2017). Recently, causal methods model and remove confounding effects (Wang et al., 2021c; Zhang et al., 2021) through causal analyses. However, these macroscopic approaches address symptoms rather than root causes (Wei et al., 2021). They fail to answer a fundamental question: how does BPR optimization geometrically distort the latent representation space to systematically favor popular items? Understanding this geometric mechanism is crucial for developing principled solutions that address bias at its source.
@@ 2.2.1. Matrix Factorization (MF)
Matrix Factorization (Rendle et al., 2009) is a foundational collaborative filtering model. It directly learns the user and item embeddings, $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$, from the interaction data. The preference score $\hat{y}_{ui}$ that user $u$ is predicted to have for item $i$ is calculated as the inner product of their respective embeddings:
@@ 2.2.2. LightGCN
LightGCN (He et al., 2020) is a state-of-the-art GNN-based model that simplifies the design of traditional Graph Convolutional Networks (GCNs) for recommendation. LightGCN begins with initial, learnable embeddings $\mathbf{E}^{(0)}=[\mathbf{e}_{u_{1}}^{(0)},...,\mathbf{e}_{i_{1}}^{(0)},...]^{T}$ and refines them through a linear propagation mechanism on the user-item interaction graph. At each layer $k$, the embeddings are aggregated from their neighbors as follows:
@@ 2.3. Bayesian Pairwise Ranking (BPR) Loss
The Bayesian Pairwise Ranking (BPR) loss is the standard optimization objective for models trained on implicit feedback datasets (Rendle et al., 2009). It is built on a pairwise learning assumption: for a given user, an observed item should be ranked higher than all unobserved items. To operationalize this, BPR constructs training instances as triplets $(u,i,j)$, where $u$ is a user, $i$ is an item they have interacted with (a positive item, $(u,i)\in\mathcal{R}$), and $j$ is an item they have not interacted with (a negative item, $(u,j)\notin\mathcal{R}$). The objective is to maximize the score difference between the positive and negative items, $\hat{y}_{ui}-\hat{y}_{uj}$. The BPR loss function is formulated as:
@@ 5.1.2. Baselines
  • Backbone Models: We show the general applicability of DDC by applying it to five representative models: MF (Rendle et al., 2009), LightGCN (He et al., 2020), DGCF (Wang et al., 2020b), NCL (Lin et al., 2022), and LightCCF (Zhang et al., 2025).
  • Debiasing Methods: We compare DDC with seven state-of-the-art debiasing methods: IPS (Joachims et al., 2017), DICE (Zheng et al., 2021), MACR (Wei et al., 2021), PC (Zhu et al., 2021b), PAAC (Cai et al., 2024), DCCL (Zhao et al., 2023), and TPAB (Yoo et al., 2025). For a fair comparison, these methods are applied to two backbone models, MF and LightGCN.
@@ 5.1.4. Implementation Details
To ensure a fair comparison, we set the training batch size to 8192 and implement all methods under the same framework, RecBole (Zhao et al., 2021, 2022; Xu et al., 2023). For all models, including backbones and debiasing baselines, we performed a grid search to find optimal hyperparameters. Crucially, we did not set an upper limit on epochs; all models were trained until convergence, determined by an early stopping strategy where training terminates if validation performance (MRR@10) does not improve for 50 consecutive epochs. The model achieving the best validation performance was then evaluated on the test set. The embedding dimension is set to $d=64$. For our DDC method, the only significant hyperparameter is the proportion of a user’s interacted items used to construct their personalized preference direction $\mathbf{e}_{pref}$. While careful tuning of this proportion can yield further gains, for the main experiments in Tables 2 and 3, we uniformly set this value to 30% to demonstrate robust performance without extensive hyperparameter tuning.
@@ 6. Related Work
Macroscopic Debiasing Strategies. Early approaches to mitigating popularity bias treat reommendation models as black boxes, intervening at the data or objective function level. For example, reweighting methods, like inversion propensity scoring (IPS) (Bottou et al., 2013; Chen et al., 2021; Zhu et al., 2021b), down-weight popular item interactions to amplify signals from niche items, but can suffer from high variance and oversimplification (Joachims et al., 2017; Schnabel et al., 2016). Regularization techniques add penalties to the objective function to discourage correlation with item popularity, often at the cost of overall recommendation accuracy (Wasilewski and Hurley, 2016; Abdollahpouri et al., 2017; Zhu et al., 2021b; Loveland et al., 2025). Critically, these macroscopic strategies address the symptoms of popularity bias without dissecting its underlying structural impact on the model’s learned representations.
Causal and Disentangled Approaches. More principled solutions leverage causal inference and representation disentanglement. Causal approaches use graphs to model and remove popularity’s confounding influence, often via counterfactual inference (Wang et al., 2021c; Zhang et al., 2021; Bonner and Vasile, 2018). Similarly, disentanglement methods aim to learn separate latent factors for genuine interest versus conformity to popular trends (Zheng et al., 2021), with the goal of isolating a ‘pure’ preference representation for prediction. While sophisticated, these methods often rely on strong causal assumptions or complex training schemes (Wei et al., 2021). Crucially, they hypothesize a popularity component but fail to explain how it is structurally and geometrically encoded by standard optimization like BPR—a foundational gap our work directly addresses.
Representation Space Optimization. Contrastive learning (CL) has emerged as a powerful paradigm to improve embedding quality in GNN-based models (Wu et al., 2021; Yu et al., 2022). CL-based methods like SGL (Wu et al., 2021) and SimGCL (Yu et al., 2022) employ a contrastive loss to foster a more uniform distribution of embeddings, which can indirectly alleviate popularity bias by preventing popular items from dominating the latent space (Wang and Isola, 2020; Wang et al., 2022). However, these approaches pursue uniformity as a general objective for representation quality; they do not directly address the specific, systematic geometric distortion that popularity imprints on the embedding space. Therefore, while beneficial, this pursuit of general uniformity is an indirect and incomplete solution to the specific geometric problem of popularity bias. Our work differs by identifying the principal “popularity direction” and then applying a targeted, asymmetric correction, fundamentally realigning the space to disentangle preference from popularity.